The primitive of the partial differential equation
√p + √q = 2x , is 1/6×(2x-a)^3+a^2y+b
True or false with full explanation
1
Expert's answer
2020-11-08T18:14:22-0500

"\displaystyle\textsf{True}\\


\sqrt{p} + \sqrt{q} = 2x\\


\sqrt{q} = 2x - \sqrt{p} = a\\


\sqrt{q} = a, q = a^2\\

2x - \sqrt{p} = a\\

\sqrt{p} = 2x - a\\

p = (2x - a)^2\\


\mathrm{d}z = p\,\mathrm{d}x + q\,\mathrm{d}y\\


\mathrm{d}z = (2x - a)^2\,\mathrm{d}x + a^2\,\mathrm{d}y\\

\int \,\mathrm{d}z = \int (2x - a)^2\,\mathrm{d}x + \int a^2\,\mathrm{d}y\\


z = \frac{(2x - a)^3}{6} + a^2y + b\\

\textsf{Therefore, it is true.}"